OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer. ln (x - 6) + ln (x + 1) = ln (x - 15)

OpenStudy (turingtest):

first apply the log rule\[\log a+\log b=\log(ab)\]